Backlinks to LabstaffDowntime in LabstaffWeb Web (Search all webs)

Results from LabstaffWeb web retrieved at 22:15 (GMT)

Mail server was not allowing people to send mail. Filesystem had filled with mimedefang files.
NEW - 24 Mar 2010 - 14:55 by lindss2@LAB.CS.RPI.EDU
March 9, 2010 * EPO to be tested during scheduled March outage * system tested and running .StevenLindsey 09 Mar 2010
r3 - 17 Mar 2010 - 14:25 by lindss2@LAB.CS.RPI.EDU
We unexpectedly lost power around 2pm in CII 3226. It appears that everything came back ok except for the chillers, but a call is in, and we're not expecting to ...
NEW - 24 Feb 2010 - 19:35 by lindss2@LAB.CS.RPI.EDU
CII MRI cluster shut down due to cooling problems 2010 05 02. Message from HVAC: Saturday morning, one chiller went in alarm and tripped due to extended surge, wh...
r2 - 03 May 2010 - 17:59 by lindss2@LAB.CS.RPI.EDU
Extensive outages caused by a failing console server that sent a break signal to all of the attached servers when it crashed. * The console server is experie...
NEW - 27 May 2010 - 14:42 by lindss2@LAB.CS.RPI.EDU
Remote access machines are being relocated for cooling purposes. If you get disconnected, simply reconnect to and you should get a machine whic...
NEW - 28 May 2010 - 14:29 by lindss2@LAB.CS.RPI.EDU
On June 3, 2010, we will be retiring the machines in Amos Eaton 215 which are hosted off of This does not affect the Semi Annual Review of Student...
r3 - 27 May 2010 - 18:09 by armstj2@LAB.CS.RPI.EDU
Location Wednesday 5/29 Thursday 5/30 Friday 5/31 Monday 6/3 Tuesday 6/4 Wednesday 6/5 Thursday 6/6 Friday 6/7 Amos Eaton (1) (1) (1) 5p...
r4 - 06 Jun 2013 - 18:51 by lindss2@LAB.CS.RPI.EDU
Location Jun 1 Jun 2 Jun 3 Jun 4 Jun 5 AE (a) (a) (a) CCNI 7am 7pm (b) 7am 7pm (b) CII 5pm 2am (c) 5pm 7pm (...
We will be transitioning the last of the networks in Amos Eaton to new routers July 12th at 10am. It will affect subnets 1, 3, 6, 8, 28, and 29. It took a whil...
r3 - 12 Jul 2011 - 17:54 by lindss2@LAB.CS.RPI.EDU
Starting a little before 3pm today, there have been authentication issues interfering with wireless networking and sending email from many clients. Sending email...
Lally and Amos Eaton will lose power for two weekends to allow the replacement of a faulty electrical feeder: 8am Friday Aug 8 4:30pm Sunday Aug 10th 8am Friday...
The CS email server will be taken down permanently in the near future, previously scheduled for October 5, 2012. This does not affect your or @alum.rpi....
r3 - 04 Oct 2012 - 18:30 by lindss2@LAB.CS.RPI.EDU
2010/09/24 CII Router Maintenance Due to scheduled maintenance, the main network router in the CII/Low Building will be down between 7:00am and 7:30am on Friday, ...
NEW - 21 Sep 2010 - 18:56 by armstj2@LAB.CS.RPI.EDU
2018 * January 2018 Outages 2017 * January 2017 Outages * May 2017 Outages 2016 * January 2016 Outages * May 2016 Outages * October 2016 Outages...
This year's electrical shutdowns are scheduled as follows: Location Fri Jun 1 Mon Jun 4 Tue Jun 5 Wed Jun 6 Thur Jun 7 Amos Eaton 5 6pm 5 6p...
r4 - 01 Jun 2012 - 14:43 by lindss2@LAB.CS.RPI.EDU
Building Time Duration Amos Eaton Monday, Tuesday, Thursday: 6am 7am 0 15 minutes CII Monday, Tuesday, Thursday: 6am 7am 5 60 minutes Lally ...
r4 - 09 Jun 2011 - 15:50 by lindss2@LAB.CS.RPI.EDU
This evening, Tuesday November 1st, beginning at approximately 5pm, maintenance to the Exchange Mail Infrastructure will begin. Maintenance may take up to 3 hrs a...
NEW - 01 Nov 2011 - 14:19 by lindss2@LAB.CS.RPI.EDU
SSL certificates expired enmasse on 2010/09/24 @ 12:18 * The problem is known and actively being worked on. Workaround * In the warning dialog box displaye...
r2 - 25 Sep 2010 - 00:05 by armstj2@LAB.CS.RPI.EDU
DotCIO will be implementing the following change: As previously announced, this summer we will start blocking access to DNS servers that are not run by RPI. The f...
r2 - 13 Jun 2012 - 20:35 by lindss2@LAB.CS.RPI.EDU
If you log in and see a message along the lines of: * Could not chdir to home directory /cs/${your userid}: Permission denied Home directories are being resto...
r13 - 20 Oct 2010 - 12:27 by lindss2@LAB.CS.RPI.EDU
The HVAC will be shut down in the MRC from Thursday, January 5th at 8am to Friday, January 6th at 4:30pm. The fans may come on for the period between midnight an...
NEW - 04 Jan 2012 - 14:52 by lindss2@LAB.CS.RPI.EDU
LOCATION Mon 2 * *Tue 3 Wed 4 impact Amos Eaton 5pm 7pm 5pm 11pm 5pm 11pm normal power outage CCNI 7am 7pm 7am 7pm sporadic emerge...
NEW - 17 Sep 2012 - 13:09 by lindss2@LAB.CS.RPI.EDU
There will be a change in the way the campus handles MX records for the CS domain. It has the potential to affect email delivery from off campus senders, but we ...
r4 - 14 Mar 2012 - 15:58 by lindss2@LAB.CS.RPI.EDU
There is a planned outage of the campus exchange email server from 5pm 6pm on Wednesday, March 13th. It is recommended (but not required) that you exit your mai...
NEW - 06 Mar 2013 - 13:39 by lindss2@LAB.CS.RPI.EDU
There will be a cooling shutdown in the MRC lasting from 8am to 4pm. We will be shutting down the grid machines named after states and research machines hosted i...
NEW - 07 Jan 2011 - 15:40 by lindss2@LAB.CS.RPI.EDU
Scheduled outages affecting CS will be as follows: Building Mon Aug 9 Tue Aug 10 Wed Aug 11 Amos Eaton 3:30pm 7:30pm 3:30pm Thursday morning ...
r3 - 06 Aug 2010 - 20:05 by lindss2@LAB.CS.RPI.EDU
I migrated our database from our FreeBSD server to an Ubuntu server on Tuesday, January 15. There was a brief outage from 9am to about 9:30am while the current d...
r2 - 15 Jan 2013 - 14:37 by lindss2@LAB.CS.RPI.EDU
There will be brief interruptions as ports are switched from our closet to the VCC closet beginning around 10am. There should be no change in the way things work ...
r4 - 15 Aug 2011 - 15:08 by lindss2@LAB.CS.RPI.EDU
Mathworks, Inc. has modified the licensing structure for Matlab. Going forward the faculty/staff license will be separate from the student license. This distincti...
NEW - 03 Aug 2012 - 13:23 by lindss2@LAB.CS.RPI.EDU
There will be a brief filesystem interruption as we re load mountd to update filesystem mounts.
NEW - 04 Mar 2011 - 15:33 by lindss2@LAB.CS.RPI.EDU
We have moved all home directories and projects onto a network appliance(NetApp). This gives us greater reliability of storage, lower maintenance, and far simple...
r4 - 26 Aug 2011 - 14:23 by lindss2@LAB.CS.RPI.EDU
There will be a number of power outages, primarily on the weekends to allow for replacing the electrical feeder. We recommend shutting down your machines ahead of...
Projects server not exporting projects, fixed
NEW - 04 Jun 2010 - 16:13 by lindss2@LAB.CS.RPI.EDU
Server for windows printing has unexpectedly powered down and was offline (2nd time). Powered back on and will be migrated onto new hardware.
r2 - 10 Jun 2010 - 15:39 by armstj2@LAB.CS.RPI.EDU
Home directories were temporarily unavailable due to server problems Drive failed and removed from raid array. Awaiting replacement Replacement installed and arr...
r4 - 09 Jun 2010 - 14:50 by lindss2@LAB.CS.RPI.EDU
Amos Eaton UPS failure Brief outages may occur as we switch from one UPS to the other UPS declared itself dead 6/10/2010
NEW - 10 Jun 2010 - 15:57 by lindss2@LAB.CS.RPI.EDU
Home directory server is currently unavailable Fixed around 2:55pm
r2 - 14 Jun 2010 - 19:51 by lindss2@LAB.CS.RPI.EDU
CII Air conditioning outage * The VCC paged at 11:22pm reporting that the chilled water temperature was rising quickly in the CII and that HVAC had been notifi...
NEW - 01 Jul 2010 - 03:58 by lindss2@LAB.CS.RPI.EDU
Project outage /projects were unavailable from around 5am until 9:23am when the outage was noticed
NEW - 01 Jul 2010 - 15:12 by lindss2@LAB.CS.RPI.EDU
CGI web server outage * stability problems with CGI server * both fans had died * seems to be better after replacement
NEW - 09 Jul 2010 - 15:51 by lindss2@LAB.CS.RPI.EDU
Outgoing mail is currently not being sent but is queueing up on the local machines Problem fixed 3:48pm, but it may take a while to clear unsent messages. Otherw...
r2 - 07 Jul 2010 - 20:12 by lindss2@LAB.CS.RPI.EDU
2010/07/08 * Network outage * Bad switch removed from service 2010/07/07 * Network instability reported * Network switch rebooted .JoeyArmstro...
NEW - 08 Jul 2010 - 16:24 by armstj2@LAB.CS.RPI.EDU
Extensive outage in MRC * Networking, some phones, and the video bulletin boards affected * VCC notified at 4:30 * Root cause unknown, but may have been ...
NEW - 09 Jul 2010 - 15:29 by lindss2@LAB.CS.RPI.EDU
Electrical problems in Amos Eaton machine room * UPS died and was causing machines attached to it to reboot occasionally at first, and then nearly constatntly ...
NEW - 10 Jul 2010 - 18:49 by lindss2@LAB.CS.RPI.EDU
After experiencing extreme network slowness, we tracked it down to packet loss going from the CS network to the VCC (but not the other direction). We will be rep...
r3 - 14 Jul 2010 - 20:21 by lindss2@LAB.CS.RPI.EDU
* OPS paged at 11:20pm on the 17th saying the cooling had failed and the temperature was going up. * No word from HVAC as of midnight, so cluster shut down ...
r2 - 20 Jul 2010 - 17:23 by lindss2@LAB.CS.RPI.EDU
Confirmed that mail seems to be piling up on the RCS exchange server for CS accounts since the 27th. Working with the VCC to fix the problem. Email from off cam...
NEW - 29 Jul 2010 - 17:37 by lindss2@LAB.CS.RPI.EDU
July Home Directory Problems * intermittent packet loss between CS networking and VCC * attempting to narrow down source of problem
NEW - 09 Jul 2010 - 15:58 by lindss2@LAB.CS.RPI.EDU
July Mail Server Problems * July 7 * many messages bounced 7/8/2010 due to dns problems * some local messages delayed or bounced * messages s...
r2 - 13 Jul 2010 - 13:31 by lindss2@LAB.CS.RPI.EDU
July Projects Directories Problems * July 9 * Lost contact with LDAP server rebooted to restore service * Multiple projects claiming to be 100% f...
NEW - 09 Jul 2010 - 16:14 by lindss2@LAB.CS.RPI.EDU
Home directories were unavailable due to an error with LDAP group generation.
NEW - 16 Aug 2010 - 13:34 by lindss2@LAB.CS.RPI.EDU
New samba server not mounting home directories Reverted to old samba server
NEW - 16 Aug 2010 - 15:11 by lindss2@LAB.CS.RPI.EDU
Samba/Windows home directories were unavailable after the server rebooted
NEW - 24 Aug 2010 - 18:31 by lindss2@LAB.CS.RPI.EDU
A brief outage occurred when the server rebooted unexpectedly
NEW - 24 Aug 2010 - 16:24 by lindss2@LAB.CS.RPI.EDU
Home directory server temporarily unavailable. NFS trouble and server panic during orderly shutdown
NEW - 09 Sep 2010 - 14:41 by lindss2@LAB.CS.RPI.EDU
Server went down at 3:32am and was restored to service 7:42am.
NEW - 10 Jan 2011 - 18:08 by lindss2@LAB.CS.RPI.EDU
Buildings affected included AE, Lally, Sage, MRC, and others. Did not include VCC, Folsom, or CII Duration around 30 minutes. Cause was said to be a squirrel sho...
NEW - 11 Jan 2011 - 21:12 by lindss2@LAB.CS.RPI.EDU
Current Status * Accounts restored as of 6am on 2/20/2011. Symptoms * zfs filesystem /cs home directories are mounted from has reported corruption. * H...
r27 - 20 Feb 2011 - 11:00 by lindss2@LAB.CS.RPI.EDU
* We were having trouble with the hard drives in the raid array, and the progress that had been made on restores was lost. * An older copy of web pages is now ...
r19 - 04 Mar 2011 - 22:25 by lindss2@LAB.CS.RPI.EDU
Laundry list * Clam Antivirus corrupt virus scanner DBs were found on two mailservers that affected receipt of incoming mail. Content has been replaced with...
NEW - 01 Mar 2011 - 01:24 by armstj2@LAB.CS.RPI.EDU
Server returned to service. No file loss.
r2 - 07 Mar 2011 - 21:04 by lindss2@LAB.CS.RPI.EDU
It looks like the server started having trouble around 4am but was available to most machines, and started having widespread trouble around 7am. The machine is b...
NEW - 08 Mar 2011 - 14:31 by lindss2@LAB.CS.RPI.EDU
At current rebuild rate, it won't be done until 9am 4/30/2011 at the earliest. 90% complete as of 8:28am Service restored 9:48 am
r3 - 30 Apr 2011 - 13:49 by lindss2@LAB.CS.RPI.EDU
The web server which serves home pages was down due to all of the httpd processes being stuck on disk wait. Server stalled during reboot and had to be power cycl...
r2 - 25 Jan 2012 - 17:53 by lindss2@LAB.CS.RPI.EDU
The air conditioner in Amos Eaton 118 stopped working. The unit turned off because the conensate pump was too full. After manually draining the pump and cleanin...
NEW - 01 Nov 2011 - 20:16 by lindss2@LAB.CS.RPI.EDU
The exhaust hose for Amos Eaton 118 (Cisco Academy) tore through in two places and could not be repaired. I was eventually able to install a new exhaust hose and ...
NEW - 04 Apr 2012 - 15:06 by lindss2@LAB.CS.RPI.EDU
The file server was on its way back up and was checking the filesystem when the watchdog timer expired and rebooted the machine Machine restored to service 8:07pm...
r2 - 04 Nov 2010 - 00:07 by lindss2@LAB.CS.RPI.EDU
Both network connections to the file server timed out and the machine had to be rebooted.
NEW - 17 Nov 2010 - 23:18 by lindss2@LAB.CS.RPI.EDU
Home directory server running in degraded mode. Users logged in may see degraded performance.
r2 - 11 Jun 2011 - 13:10 by lindss2@LAB.CS.RPI.EDU
The CGI server ran out of swap space during project submission and crashed and would not boot due to the IPMI log being full. Booted to BIOS and cleared out yea...
NEW - 28 Oct 2011 - 13:22 by lindss2@LAB.CS.RPI.EDU
Cooling is back, and not expected to need to be shut down in an emergency situation as the load has been rebalanced, but the electrical work is not all done, so a...
r3 - 23 Jun 2011 - 12:46 by lindss2@LAB.CS.RPI.EDU
An emergency power outage is being scheduled by the RPI Electrical shop to replace an emergency transfer switch at the CII. The outage is scheduled to beginning ...
NEW - 15 Nov 2012 - 20:31 by lindss2@LAB.CS.RPI.EDU
The following message came out warning of a brief outage Friday, November 30 starting at 6am. Please plan accordingly. DotCIO will be implementing some changes t...
NEW - 28 Nov 2012 - 17:46 by lindss2@LAB.CS.RPI.EDU
DNS service is now handled by the VCC, but short term please continue to send requests through labstaff #64; until troubleshooting is complete.
r6 - 03 Aug 2011 - 12:45 by lindss2@LAB.CS.RPI.EDU
The project server crashed twice when trying to delete a project that was interfering with backups. We will leave it until the transition to the netapp is comple...
NEW - 29 Jun 2011 - 14:16 by lindss2@LAB.CS.RPI.EDU
Due to some file corruption in the email server config, some people were not able to delete email. Problems should all have been fixed by 2am Oct 3.
NEW - 03 Oct 2012 - 13:43 by lindss2@LAB.CS.RPI.EDU
OutageGithub is blocked by the VCC at the moment because it wound up on a blacklist (Botnet Traffic Filter). They said that someone from github will need to talk t...
NEW - 28 Oct 2010 - 19:48 by lindss2@LAB.CS.RPI.EDU
The unix print server rebooted 3 times last night. Printing directly to printer not affected.
NEW - 13 Feb 2012 - 15:54 by lindss2@LAB.CS.RPI.EDU
Server mysteriously rebooted twice and won't respond now. Printing directly to printers from desktops or not affected, but printing from remote.c...
r3 - 07 May 2012 - 16:17 by lindss2@LAB.CS.RPI.EDU
The server on which home directories and the majority of our infrastructure depends is no longer able to boot properly. I finished copying the information off of...
r4 - 01 Aug 2011 - 14:01 by lindss2@LAB.CS.RPI.EDU
One of the campus LDAP servers run by CMT,, started having disk issues late last week. Saturday, it effectively died. As it turns out, a lot of...
NEW - 03 Sep 2013 - 20:31 by lindss2@LAB.CS.RPI.EDU
CS Mailing lists were not working properly. Old messages should all be sent.
r2 - 06 Mar 2012 - 18:05 by lindss2@LAB.CS.RPI.EDU
The network transition is complete. Some ports on subnet 17 had a configuration problem, but should be fixed as of 3:30pm 2/15/2012. If you are still having any ...
r3 - 15 Feb 2012 - 20:49 by lindss2@LAB.CS.RPI.EDU
OutageMRCCooling is available again providing the air conditioning continues to work and at least some of the machines should be available during the pow...
r6 - 06 Jun 2011 - 15:49 by lindss2@LAB.CS.RPI.EDU
Cooling should be stable, although another outage to finish fixing the electrical infrastructure may still be needed in the future. If is not av...
r7 - 23 Jun 2011 - 12:48 by lindss2@LAB.CS.RPI.EDU
Mail server was dependent on several other servers, and we were hoping to migrate away from it before problems hit. Unfortunately that was not possible, so I con...
r7 - 01 Aug 2011 - 15:47 by lindss2@LAB.CS.RPI.EDU
It appears that email sent to (CS) mailing lists started bouncing January 24th or 25th. Regular email does not appear to be affected. Missed messages will still...
r2 - 27 Jan 2013 - 10:52 by lindss2@LAB.CS.RPI.EDU
Mailing lists broken as of November 11th at 4:59 pm. Problem seems to have come from nscd caching bad data. Still trying to fix problems that stemmed from it. Ma...
r2 - 12 Nov 2010 - 20:01 by lindss2@LAB.CS.RPI.EDU
Due to an air conditioning outage in the MRC, we have shut down all of the machines in the MRC machine room except for the networking switches, which we hoped wou...
r11 - 07 Oct 2011 - 16:02 by lindss2@LAB.CS.RPI.EDU
Due to a misunderstanding, the partition on the NetApp that most of the personal web directories were mounted on was was believed to be inactive and destroyed Jul...
r13 - 19 Jul 2011 - 16:13 by lindss2@LAB.CS.RPI.EDU
Those users not transitioned to exchange and CS mailing lists not working due to a mistake in DNS. Fixed around 1:30pm. It turned out that problems from the chang...
r3 - 01 Dec 2011 - 21:48 by lindss2@LAB.CS.RPI.EDU
The server reports that the user is not known when trying to send from the exchange server to a non exchange email address even when using the web client. It appe...
NEW - 25 Sep 2012 - 18:54 by lindss2@LAB.CS.RPI.EDU
Some users seem to have corrupt entries that allow them to get their mail, but not log in to respite or the VPN. Changing the password sometimes fixes the proble...
NEW - 23 Jul 2012 - 18:34 by lindss2@LAB.CS.RPI.EDU
The mailing list server went down and was up as of 12:25pm.
NEW - 16 May 2012 - 16:25 by lindss2@LAB.CS.RPI.EDU
The subversion and trac server appeared to have lost a power supply. Since it was going to take over 2 weeks to fix it, we transferred the service to another mac...
r4 - 22 Jan 2014 - 20:40 by lindss2@LAB.CS.RPI.EDU
While trying to restore an old web page for a user, the filesystem mounts got messed up. Everything should be back to normal as of 10:20pm 2/15/2012. Outage due ...
r2 - 29 Feb 2012 - 16:03 by lindss2@LAB.CS.RPI.EDU
FreeBSD workstations depended on barak and will need to be re installed with Ubuntu when time permits (not for a while) Solaris workstations will need to be conve...
r3 - 28 Oct 2011 - 13:23 by lindss2@LAB.CS.RPI.EDU
Physical Facilities said that the MRC is in a shutdown and would be for at least 2 days. Shutting down machines in the MRC machine room to prevent heat damage. ...
r7 - 05 Jan 2011 - 14:37 by lindss2@LAB.CS.RPI.EDU
URLs * Email Help Configuration * Webmail Problems 2010 06 04 * Intermittent problems have been reported. * remote clients may have problems conn...
r2 - 07 Jun 2010 - 09:37 by armstj2@LAB.CS.RPI.EDU
We are expiring passwords for all users that have not changed their passwords since January 2011. You will be required to log in to (or one of ...
r6 - 07 Jun 2011 - 20:40 by lindss2@LAB.CS.RPI.EDU
There is a power outage scheduled for Amos Eaton, Lally, and the Green building from 1pm on August 1st to 3am on August 2nd. Email and will not...
r6 - 02 Aug 2012 - 10:53 by lindss2@LAB.CS.RPI.EDU
5 7pm Remove breaker for repair. This is phase 1 of 3. Two additional outages will be required to complete this work. The next two dates are TBD. See also http...
NEW - 03 Jul 2012 - 14:04 by lindss2@LAB.CS.RPI.EDU
Power will be out for portions of 4:30pm to 6:30pm on Tuesday, August 28th. Please see for details.
NEW - 21 Aug 2012 - 12:57 by lindss2@LAB.CS.RPI.EDU
Building 5/31 6/1 6/2 6/3 Amos Eaton 5pm 7pm 5pm 7pm 5pm 7pm CII 1pm 2am 5pm 7pm 5pm 7pm Lally 5pm 7pm 5pm 7pm 5pm 7pm ...
location July 3 July 5 July 6 July 7 *July 10 * Amos Eaton N/A 4:30pm 9:30pm (a) 4:30pm 8:30pm (a) 4:30 6:30pm (b) all day (c) CII all ...
Reporting a problem Before reporting a problem, please check the downtime page to see if the problem has already been reported. Problems with email should be repo...
r5 - 16 Aug 2012 - 13:33 by lindss2@LAB.CS.RPI.EDU
Last Published Publisher JoeyArmstrong Date 02 Jan 2008 19:48 {PublishContrib}{Dir} /local/TWiki/published/ {PublishContrib}{URL}
r30 - 08 Jan 2010 - 00:16 by armstj2@LAB.CS.RPI.EDU
RPICS wireless service will be discontinued June 30th, 2010. You should transition to 802.1x wireless ahead of time and let us know of any difficulties you have....
r3 - 29 Jun 2010 - 15:49 by lindss2@LAB.CS.RPI.EDU
/projects will be temporarily unavailable around 1am for system maintenance
NEW - 07 Jun 2010 - 16:24 by lindss2@LAB.CS.RPI.EDU
We transitioned to having the VCC provide dns on Wednesday, August 3rd at 8am due to technical problems. If you manually configure your DNS settings, it should be...
r4 - 03 Aug 2011 - 14:07 by lindss2@LAB.CS.RPI.EDU
We are transitioning to email being hosted on the campus exchange server. During the transition, you can check to see if it is complete by sending email to your ...
r35 - 10 Aug 2012 - 14:51 by lindss2@LAB.CS.RPI.EDU
Many students have had their CS email migrated to being delivered to their CS accounts. Please be sure to update your Respite settings to be sure your spam filte...
NEW - 26 Mar 2012 - 19:27 by lindss2@LAB.CS.RPI.EDU
The CS mailing lists transitioned to the Sympa server in the VCC, and are accessible from In brief: * The new server has a lot more ...
r8 - 30 Aug 2012 - 16:32 by lindss2@LAB.CS.RPI.EDU
For all regular requests, send email to labstaff #64; If you need to page labstaff, faculty and staff can email labstaff pager #64; Please ...
Important Note * Do not fall for scams that try to gather passwords or personal information. (Click here for tips on avoiding them) * Alerts and Downtime ...
Statistics for LabstaffWeb Web Month: Topic views: Topic saves: File uploads: Most popular topic views: Top contributors for topic save and uploa...
February 17, 2010 * zeus having hardware problems causing instability * removed 10G networking card to restore stability .StevenLindsey 18 Feb 2010
NEW - 18 Feb 2010 - 20:22 by lindss2@LAB.CS.RPI.EDU
Zeus continues to have problems with one of the partitions. We hope to have it back up on the 5th or 6th after moving data off of the affected partition. .Stev...
NEW - 05 Mar 2010 - 04:29 by lindss2@LAB.CS.RPI.EDU
Number of topics: 120
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback